OpenStudy (anonymous):

Which of the following describes the translation of the graph of y = x2 to obtain the graph of y = -x2 + 3? reflect over the x-axis and shift right 3 reflect over the x-axis and shift up 3 reflect over the y-axis and shift down 3

OpenStudy (jdoe0001):

\(\textit{function transformations} \\ \quad \\ \begin{array}{llll} \begin{array}{llll} shrink\ or\\ expand\\ by\ {\color{purple}{ A}}\cdot {\color{blue}{ B}}\end{array} \qquad \begin{array}{llll} vertical\\ shift\\ by \ {\color{green}{ D}} \end{array} \begin{array}{llll}{\color{green}{ D}} > 0& Up\\ {\color{green}{ D}} < 0 & Down\end{array} \\ \qquad \downarrow\qquad\qquad\qquad \downarrow\\ y = {\color{purple}{ A}} ( {\color{blue}{ B}}x + {\color{red}{ C}} ) + {\color{green}{ D}}\\ \qquad\qquad\uparrow \\ \qquad\begin{array}{llll} horizontal\\ shift\\ by \ \frac{{\color{red}{ C}}}{{\color{blue}{ B}}}\end{array} \begin{array}{llll}\frac{{\color{red}{ C}}}{{\color{blue}{ B}}} > 0 & to \ left\\ \frac{{\color{red}{ C}}}{{\color{blue}{ B}}} < 0& to \ right\end{array} \end{array}\) any ideas? bear in mind that, a minus in front of the leading term, flips the equation upside down